The Warren W. Fisher Computer Corporation purchases 10,000 transistors each year as components in minicomputers. The unit cost of each transistor is $10, and the cost of carrying one transistor in inventory for a year is 25% of the unit cost. Ordering cost is $20 per order. If the lead time is 4 days, at which inventory level should the manager request a delivery (assume basic EOQ is used and the company operates a total of 250 workdays per year)?
Answer #1

Annual demand (D) = 10000 units

Number of days per year = 250 days

Average daily demand(d) = D/number of days per year = 10000/250 = 40

Lead time (L) = 4 days

Reorder point = d x L = 40 × 4 = 160 units

So at an inventory level of 160 units the manager should request a delivery